How many moles of Cl2 can be produced from 5.60 mol HCl in reaction
HCl + O2 → Cl2 + H2O (unbalanced) (3 significant. final answer)
1
Expert's answer
2020-06-02T14:08:25-0400

The balanced reaction is:

4HCl + O2 → 2Cl2 + 2H2O.

According to the reaction, n(Cl2) = 1/2 n(HCl) = 5.60/2 = 2.80 moles.

Answer: 2.80 moles